OpenStudy (anonymous):

The function represents a composition of two functions. Identify the two functions for h(x)=(3x^2+x+1)^3

OpenStudy (anonymous):

f(x) = x^3 and g(x) = 3x^2 + x + 1. Then f(g(x)) = h(x).